find the measures of an angle and its complement if one angle measures 18 degrees more than the other

Answer by stanbon(75887)

find the measures of an angle and its complement if one angle measures 18 degrees more than the other
------
1st angle : x
2nd angle: 90-x
-----
90-x = x+18
2x = 72
x = 36 degrees
90-x = 54 degrees
=======================
Cheers,
Stan H.